OpenStudy (anonymous):

Find two numbers whose difference is 150 and whose product is a minimum.

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hi!

OpenStudy (misty1212):

lets call one of the numbers \(x\) then the other must be \(150+x\) since there difference is \(150+x-x=150\)

OpenStudy (misty1212):

the product is therefore \[P(x)=x(150+x)\] and you want to minimize that one

OpenStudy (misty1212):

\[P(x)=150x+x^2\] is a parabola that opens up the minimum is at the vertex the first coordinate of the vertex of \(y=ax^2+bx+c\) is \(-\frac{b}{2a}\) which in your example is \[\frac{150}{2}=-75\]

OpenStudy (misty1212):

that makes the other one \(150-75=75\) and those are your two numbers

OpenStudy (anonymous):

So you don't need to take the derivative of P(x)=150x+x^2?

OpenStudy (misty1212):

lol no not unless you are a slave to calculus no one needs calculus to find the vertex of a parabola but you will get the same answer if you do it that way

OpenStudy (misty1212):

the derivative is \(150-2x\) set it equal to zero and you still get \(-75\)

OpenStudy (misty1212):

oops i meant the derivative is \(150+2x\) doe

OpenStudy (anonymous):

Okay, Thank you! :) @misty1212

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]